OpenStudy (anonymous):

integration of y(1+y)/(1-y)

OpenStudy (shubhamsrg):

y(1+y)/(1-y) = - (1-y -1)(1+y)/(1-y) = -(1 - 1/(1-y))*(1+y) = -(1+y) + (1+y)/(1-y) hope this simplifies ? i mean hope this helps .